given the equation: y-2= -|x-3| + 1
A. sketch the graph of the equatiion. label three points on graph
b. find the x and y intercept of the graph

Answer by josgarithmetic(39618)

Simplify first and then look at two conditions.

y=-abs(x-3)+1+2
SIMPIFLIED: y=-abs(x-3)+3

x-3<0. Corresponds to x<3.
y=-(3-x)+3
y=x-3+3
y=x

x-3>=0. Corresponds to x>=3.
y=-(x-3)+3
y=-x+3+3
y=-x+6

WHAT TO GRAPH:
------------------------------
y=x________for x<3
y=-x+6_____for x>=3
------------------------------